# From Sequence to Structure: Uncovering Substructure Reasoning in Transformers

### v1

@@ 1 Introduction
In this work, we explore how Transformers tackle the broader challenge of substructure understanding, with a particular emphasis on the task of substructure extraction. Building on the use of LLMs for substructure extraction [6, 3],  Figure 1a) illustrates the overall process with Transformers: Transformer-based models receive a query prompt, which is composed of a textual graph representation and a question prompt as input. Then, they identify the relevant substructure, and generates an answer for the given graph.
To investigate how Transformers derive answers from input tokens, in Section 3 we conduct empirical and theoretical analyses of train-from-scratch Transformers, focusing on the internal Transformer mechanisms and input queries, as shown in  Figure 1b). To understand the internal behavior of the model, we introduce a new perspective, Induced Substructure Filtration (ISF), which suggests that Transformers perform a layer-wise node aggregation process to detect substructures.
To verify the reliability of our interpretation modules, we demonstrate that our approach also applies to understanding LLM behavior and argue that it can inform the development of future methods. In Section 4, we show that our explanation for Transformers aligns with the behaviors observed in LLMs, particularly in how they tackle various textual graph representations and perform graph extraction tasks. Furthermore, in Section 5, we explore the potential of Transformers in graph understanding, building on insights from our interpretation modules. We argue that it is reasonable to extend Transformers to handle attributed graphs, such as molecular graphs. Finally, we introduce the Thinking-In-Substructure framework, which enhances Transformers’ capabilities in complex graph reasoning tasks. In summary, we offer a new perspective on how Transformers understand graph structures from sequential inputs. Our key contributions are summarized as follows:
@@ 2.1 Problem formulations
Although we prompt LLMs to interpret natural language sentences to answer substructure-related questions, we simplify the process by converting these sentences into symbolic tokens. This enables us to analyze and train Transformers from scratch. The simplified process is illustrated in Figure 1 c). Given a graph $G=\{V,E\}$ and a question prompt $T$, we encode them using $\mathrm{Encoder_{G}}$ and $\mathrm{Encoder_{T}}$ respectively to obtain simplified sentence sequences. These are then concatenated, with the encoded question placed after the graph representation, forming the input query $Q$. The objective is to extract the set of isomorphic subgraphs $\hat{G}=\{g_{1},g_{2},\cdots,g_{s}\}$, where each $g_{i}$ represents an instance of the desired substructure within the input graph. The overall process is defined as: $\mathrm{Encoder_{A}}(\hat{G})=\mathrm{Transformers}(Q)=\mathrm{Transformers}(\mathrm{Encoder_{G}}(G),\mathrm{Encoder_{T}}(T))$. We introduce each encoder of this framework in the following paragraphs.
@@
The details of the definations are in the Definition A.1 and Definition A.2 in Section A.1.
@@ 2.2 Experiment settings
We train Transformer models using the same architecture as GPT-2 but in a lightweight version, with only 384 hidden dimensions and a small number of layers depending on the tasks. The details for each task are shown in the Appendix B. During training, the model is optimized only to predict $\mathsf{ANS}(\hat{G})$. For evaluation, we use accuracy as the metric. A predicted answer is considered correct only if the $\mathsf{ANS}(\hat{G})$ is exactly the same with the ground truth.
@@ 3 Interpretations for Substructure Extraction in Transformers
In this section, we present our insights into how Transformers perform substructure understanding. We focus on two main aspects: the internal mechanisms of Transformers in solving the substructure extraction task, discussed in Section 3.1, and the impact of input query formulation on extraction performance Section 3.2.
@@ 3.1 Induced Substructure Filtration in Transformer
In this subsection, we introduce how Transformers solve the substructure extraction task. First, we show that Transformers can extract substructures of diverse shapes, as detailed in Section 3.1.1. We then analyze the underlying mechanism and propose the ISF process in Section 3.1.2. Finally, we demonstrate how ISF generalizes to cases with multiple substructures of varying numbers and shapes in Section 3.1.3.
@@ 3.1.1 Single substructure extraction
We begin by evaluating whether Transformers can extract a single specific target substructure from a given graph, which may vary in both scale and shape. The selected substructures contain 3 to 5 nodes and 3 to 6 edges. We also investigate the effects of dataset size and the number of Transformer layers. To this end, we vary the training set size from 100K to 400K and the number of Transformer layers from 2 to 5. For each substructure, we evaluate the extraction accuracy on 30K test graphs, averaging results over three runs. Table 1 shows the results for various substructure extraction tasks.
@@
# Training # Layer Triangle Path Square Diagnoal T_triangle F_Triangle Diamond Pentagon House
@@ 3.1.2 Induced Subgraph Filtration
We visualize token embeddings to better understand how Transformers extract substructures. Since decoder-only Transformers process input left to right [2], we use the final token embeddings to reflect their graph understanding. We apply t-SNE to project these embeddings into a 2D space, labeling each graph by its substructure answer, which is represented by the node IDs as illustrated in $\mathsf{ANS}(\hat{G})$. As an example, we use the square substructure extraction task, shown in Figure 4.
Figure 2: Visualization across 4 layers. We show the node ID distributions of target substructures.
@@
We further provide a theoretical framework for this process by introducing filtrations to formalize progressive substructure extraction. Using the square extraction visualization as an example, we model it as a 4-node, 3-filtration process, indicating that the target substructure contains 4 nodes and requires 3 filtration steps, each corresponding to an induced subgraph. More generally, we define this framework as a $k$-Node $m$-Filtration, referred to as Induced Subgraph Filtration, as defined in Definition 3.1.
@@
Figure 4 illustrates the concept defined in Definition 3.1. We model the extracted substructure as $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k$, and represent the extraction process using an $m$-Filtration, where $m$ denotes the number of gathering operations required for the Transformers to identify the substructure.
@@
For graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), the subgraph isomorphism indicator tensor $\mathcal{T}(G,G^{\prime})$ is $k$-dimensional ($n\times\dots\times n$) where its entry for an ordered $k$-tuple of vertices $(v_{j_{1}},\dots,v_{j_{k}})$ from $V$ is $1$ if these vertices induce a subgraph isomorphic to $G^{\prime}$ (via a predefined mapping $v^{\prime}_{p}\mapsto v_{j_{p}}$ for $p=1,\dots,k$), and $\mathcal{T}_{j_{1},\dots,j_{k}}\leq 0$ otherwise (see Definition A.5 for details).
Theorem 3.3 (proof in Section A.3) shows that Transformers can progressively compute $\mathcal{T}(G,G^{\prime})$ for each substructure along the filtration. $O(n^{k})$ is the hidden dimension needed for the Transformer to check all $k$-node subgraphs in an $n$-node graph.
@@
Furthermore, the Transformer extracts the unique instance of $G^{\prime}$, meaning the answer is uniquely determined by the given graph representation and question prompt, and $\mathcal{T}(G,G^{\prime})$ contains exactly one entry equal to 1. This leads to 3.4 and Theorem 3.5. With this condition, the substructure extraction task is solvable for transformers.
Assumption 3.4 (Unique Instance).
@@
Under 3.4, for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
@@
Theorems 3.3 and 3.5 hold for various input graph representations (e.g., adjacency lists $\mathsf{AL}(G)$ or edge lists $\mathsf{EL}(G)$), as formally defined in Definitions A.1 and A.2.
@@ 3.1.3 Simultaneous detection of multiple substructures
As graphs often contain multiple and diverse substructures, we further demonstrate how the ISF process adapts to such scenarios. Specifically, we evaluate whether a 4-layer Transformer can accurately detect both repeated and differently shaped substructures. To this end, we design the Multi-Num and Multi-Shape evaluation settings, with the pipeline illustrated in Figure 4.
Multi-Num
As shown in Figure 4 a), the Multi-Num task involves training and testing on graph sets where each sample may contain multiple target substructures—up to five in total. This task evaluates whether Transformers can successfully extract all target substructures within a single graph. Both triangles and square detections can achieve over 85% accuracy, as shown in Figure 8. The number of substructures has little impact on the Transformers’ ability. They can identify multiple patterns at once. We further analyze examples with two substructures and find that answers are still often determined before the final generation step (Figure 8).
Multi-Shape
As illustrated in Figure 4 b), the Multi-Shape task involves training and testing on graph sets where each sample may contain multiple substructures of different shapes. This task evaluates whether Transformers can identify diverse substructure types within a single graph. We create this task by mixing single-shape training data and dividing it into four groups (details in Section B.2), as shown in Figure 8. Each substructure can be extracted independently, guided by its specific question prompt. Therefore, in visualization, we take the embedding of the last input query token rather than the last graph representation token. From Figure 8, we find that Transformers often identify simpler substructures, like triangles, in layer 3, instead of delaying all predictions to the final layer (the Transformer has 4 layers in this setting).
@@
Both subtasks perform well with just a 4-layer Transformer, suggesting that simultaneous ISF processing allows effective substructure reasoning even at shallow depths. We also provide theoretical support for Multi-Num (Theorem 3.7) and Multi-Shape (Theorem 3.8) extraction, with proofs in Section A.3.
Theorem 3.7 (Expressiveness for Multi-Num Extraction).
For directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output all $k$-tuples of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
Theorem 3.8 (Expressiveness for Multi-Shape Extraction).
For a directed graph $G=(V,E)$ ($|V|=n$) and any target subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k^{\prime}\leq k$ satisfying 3.4, a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k^{\prime}$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k^{\prime}}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k^{\prime}}}=1$.
Remark 3.9.
Theorems 3.7 and 3.8 also hold for various input graph representations (e.g., adjacency lists $\mathsf{AL}(G)$ or edge lists $\mathsf{EL}(G)$), as formally defined in Definitions A.1 and A.2.
@@ 3.2 Impact of Input Query Formulation
As illustrated in Section 2, the input query consists of two components: a text-based graph representation and a question prompt. In Section 3.2.1 and Section 3.2.2, we discuss how Transformers perform the substructure extraction task based on these inputs.
@@ 3.2.1 Text-Based Graph Representations
We start with the comparison of different text-based graph representation methods. Specifically, we focus on the two basic methods, which are the neighborhood-based AL and the edge-based EL. To ensure controllable input lengths, we conduct a toy experiment using graphs with 4 to 8 nodes, keeping both representations within a 100-token limit. We then vary the number of Transformer layers to extract substructures (e.g., triangle, square, or pentagon) from the graph representations. Experimental details are provided in the Table 6 in Section B.1 and the results are shown in Figure 9.
@@
The experimental results indicate that both AL and EL formats allow Transformers to extract substructures from text-based graph representations. As the size of the target substructure increases, both formats require more Transformer layers to achieve more than 80% accuracy. As mentioned in Remarks 3.6 and 3.9, the theoretical results hold for both AL and EL formats. The intuition is that both formats can be transformed into the same binary adjacency matrix $A(G)$, which encodes the structure of graph $G$. This matrix is then vectorized as $\mathsf{vec}(A(G))$ for processing within the model, where we provides the details in Lemma A.6 in Section A.2.
Although we theoretically show that EL and AL are equivalent in representational power, experimental results reveal that EL requires more Transformer layers to achieve comparable performance. This may be because EL inherently requires more tokens to explicitly represent all edges, whereas AL encodes the same information with fewer tokens and additional padding, benefiting from its more compact structure. In practice, for a fully connected graph, the AL needs $2\times|V|^{2}-|V|$ tokens, while the EL needs $3\times(|V|^{2}-|V|)$ tokens. Therefore, for efficiency, we mainly adopt AL in our discussions.

### v2

@@ 1 Introduction
In this work, we explore how Transformers tackle the broader challenge of substructure understanding, with a particular emphasis on the task of substructure extraction. Building on the use of LLMs for substructure extraction [7, 3],  Figure˜1a) illustrates the overall process with Transformers: Transformer-based models receive a query prompt, which is composed of a textual graph representation and a question prompt as input. Then, they identify the relevant substructure, and generates an answer for the given graph.
To investigate how Transformers derive answers from input tokens, in Section˜3 we conduct empirical and theoretical analyses of train-from-scratch Transformers, focusing on the internal Transformer mechanisms and input queries, as shown in  Figure˜1b). To understand the internal behavior of the model, we introduce a new perspective, Induced Substructure Filtration (ISF), which suggests that Transformers perform a layer-wise node aggregation process to detect substructures.
To verify the reliability of our interpretation modules, we demonstrate that our approach also applies to understanding LLM behavior and argue that it can inform the development of future methods. In Section˜4, we show that our explanation for Transformers aligns with the behaviors observed in LLMs, particularly in how they tackle various textual graph representations and perform graph extraction tasks. Furthermore, in Section˜5, we explore the potential of Transformers in graph understanding, building on insights from our interpretation modules. We argue that it is reasonable to extend Transformers to handle attributed graphs, such as molecular graphs. Finally, we introduce the Thinking-In-Substructure framework, which enhances Transformers’ capabilities in complex graph reasoning tasks. In summary, we offer a new perspective on how Transformers understand graph structures from sequential inputs. Our key contributions are summarized as follows:
@@ 2.1 Problem formulations
Although we prompt LLMs to interpret natural language sentences to answer substructure-related questions, we simplify the process by converting these sentences into symbolic tokens. This enables us to analyze and train Transformers from scratch. The simplified process is illustrated in Figure˜1 c). Given a graph $G=\{V,E\}$ and a question prompt $T$, we encode them using $\mathrm{Encoder_{G}}$ and $\mathrm{Encoder_{T}}$ respectively to obtain simplified sentence sequences. These are then concatenated, with the encoded question placed after the graph representation, forming the input query $Q$. The objective is to extract the set of isomorphic subgraphs $\hat{G}=\{g_{1},g_{2},\cdots,g_{s}\}$, where each $g_{i}$ represents an instance of the desired substructure within the input graph. The overall process is defined as: $\mathrm{Encoder_{A}}(\hat{G})=\mathrm{Transformers}(Q)=\mathrm{Transformers}(\mathrm{Encoder_{G}}(G),\mathrm{Encoder_{T}}(T))$. We introduce each encoder of this framework in the following paragraphs.
@@
The details of the definitions are in the Definition˜D.1 and Definition˜D.2 in Section˜D.1.
@@ 2.2 Experiment settings
[…]
@@ 3 Interpretations for Substructure Extraction in Transformers
In this section, we present our insights into how Transformers perform substructure understanding. We focus on two main aspects: the internal mechanisms of Transformers in solving the substructure extraction task, discussed in Section˜3.1, and the impact of input query formulation on extraction performance Section˜3.2.
@@ 3.1 Induced Substructure Filtration in Transformer
In this subsection, we introduce how Transformers solve the substructure extraction task. First, we show that Transformers can extract substructures of diverse shapes, as detailed in Section˜3.1.1. We then analyze the underlying mechanism and propose the ISF process in Section˜3.1.2. Finally, we demonstrate how ISF generalizes to cases with multiple substructures of varying numbers and shapes in Section˜3.1.3.
@@ 3.1.1 Single substructure extraction
We begin with the Single-Shape-Single-Num case, evaluating whether Transformers can extract a specific target substructure from a given graph whose scale and shape may vary. The selected substructures contain 3 to 5 nodes and 3 to 6 edges. We also investigate the effects of dataset size and the number of Transformer layers. To this end, we vary the training set size from 100K to 400K and the number of Transformer layers from 2 to 5. For each substructure, we evaluate the extraction accuracy on 30K test graphs, averaging results over three runs. Table˜1 shows the results for various substructure extraction tasks.
@@
# Training # Layer Triangle Path Square Diagonal T_triangle F_Triangle Diamond Pentagon House
@@ 3.1.2 Induced Subgraph Filtration
[…]
Figure 2: Visualization across 4 layers. We show the node ID distributions of target substructures. Legends indicate the node IDs
@@
We further provide a theoretical framework for this process by introducing filtrations to formalize progressive substructure extraction. Using the square extraction visualization as an example, we model it as a 4-node, 3-filtration process, indicating that the target substructure contains 4 nodes and requires 3 filtration steps, each corresponding to an induced subgraph. More generally, we define this framework as a $k$-Node $m$-Filtration, referred to as Induced Subgraph Filtration, as defined in Definition˜3.1.
@@
Figure˜4 illustrates the concept defined in Definition˜3.1. We model the extracted substructure as $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k$, and represent the extraction process using an $m$-Filtration, where $m$ denotes the number of gathering operations required for the Transformers to identify the substructure.
@@
For graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), the subgraph isomorphism indicator tensor $\mathcal{T}(G,G^{\prime})$ is $k$-dimensional ($n\times\dots\times n$) where its entry for an ordered $k$-tuple of vertices $(v_{j_{1}},\dots,v_{j_{k}})$ from $V$ is $1$ if these vertices induce a subgraph isomorphic to $G^{\prime}$ (via a predefined mapping $v^{\prime}_{p}\mapsto v_{j_{p}}$ for $p=1,\dots,k$), and $\mathcal{T}_{j_{1},\dots,j_{k}}\leq 0$ otherwise (see Definition˜D.5 for details).
Theorem˜3.3 (proof in Section˜D.3) shows that Transformers can progressively compute $\mathcal{T}(G,G^{\prime})$ for each substructure along the filtration. $O(n^{k})$ is the hidden dimension needed for the Transformer to check all $k$-node subgraphs in an $n$-node graph.
@@
Furthermore, the Transformer extracts the unique instance of $G^{\prime}$, meaning the answer is uniquely determined by the given graph representation and question prompt, and $\mathcal{T}(G,G^{\prime})$ contains exactly one entry equal to 1. This leads to ˜3.4 and Theorem˜3.5. With this condition, the substructure extraction task is solvable for transformers.
Assumption 3.4 (Single-Shape-Single-Num).
@@
Under ˜3.4, for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
@@
Theorems˜3.3 and 3.5 hold for various input graph representations (e.g., adjacency lists $\mathsf{AL}(G)$ or edge lists $\mathsf{EL}(G)$), as formally defined in Definitions˜D.1 and D.2.
@@ 3.1.3 Simultaneous detection of multiple substructures
As graphs often contain multiple and diverse substructures, we further demonstrate how the ISF process adapts to such scenarios. Specifically, we evaluate whether a 4-layer Transformer can accurately detect both repeated and differently shaped substructures. To this end, we design the Single-Shape-Multi-Num and Multi-Shape-Single-Num evaluation settings, with the pipeline illustrated in Figure˜4.
Single-Shape-Multi-Num
As shown in Figure˜4 a), the Single-Shape-Multi-Num task involves training and testing on graph sets where each sample may contain multiple target substructures—up to five in total. This task evaluates whether Transformers can successfully extract all target substructures within a single graph. We define the task in Definition˜3.7.
Definition 3.7 (Single-Shape-Multi-Num Extraction).
The Single-Shape-Multi-Num extraction task requires a model to output all $k$-tuples of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ corresponding to occurrences of directed graph $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$) in $G=(V,E)$ ($|V|=n$). Formally, the objective is to output all tuples satisfying ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$, where ${\mathcal{T}}(G,G^{\prime})$ is Subgraph Isomorphism Indicator Tensor defined in Definition˜3.2.
As shown in Figure˜8, Both triangles and square detections can achieve over 85% accuracy. The number of substructures has little impact on the Transformers’ ability. They can identify multiple patterns at once. We further analyze examples with two substructures and find that answers are still often determined before the final generation step (Figure˜8). The theoretical explanation is provided in Theorem˜3.8.
Theorem 3.8 (Expressiveness for Single-Shape-Multi-Num Extraction).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension that can complete Single-Shape-Multi-Num Extraction defined in Definition˜3.7 for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$).
@@
Multi-Shape-Single-Num
As illustrated in Figure˜4 b), the Multi-Shape-Single-Num task involves training and testing on graph sets where each sample may contain multiple substructures of different shapes. This task evaluates whether Transformers can identify diverse substructure types within a single graph defined in Definition˜3.9.
Definition 3.9 (Multi-Shape-Single-Num Extraction).
The Multi-Shape-Single-Num extraction task requires a model to find all the occurrences for any directed graph $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|\leq k$) in $G=(V,E)$ ($|V|=n$) satisfying ˜3.4. Formally, the objective is to output the unique $k^{\prime}$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k^{\prime}}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k^{\prime}}}=1$, where ${\mathcal{T}}(G,G^{\prime})$ is Subgraph Isomorphism Indicator Tensor defined in Definition˜3.2.
We create this task by mixing single-shape training data and dividing it into four groups (details in Section˜E.2), as shown in Figure˜8. Each substructure can be extracted independently, guided by its specific question prompt. Therefore, in visualization, we take the embedding of the last input query token rather than the last graph representation token. From Figure˜8, we find that Transformers often identify simpler substructures, like triangles, in layer 3, instead of delaying all predictions to the final layer (the Transformer has 4 layers in this setting). We summarize the mechanism in Theorem˜3.10.
Theorem 3.10 (Expressiveness for Multi-Shape-Single-Num Extraction).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension that can complete Multi-Shape-Single-Num Extraction defined in Definition˜3.9 for a directed graph $G=(V,E)$ ($|V|=n$) and any target subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k^{\prime}\leq k$ satisfying ˜3.4.
These two properties form the foundation for understanding how decoder-only Transformers decompose complex graphs into simpler ones for substructure extraction, as discussed in Section˜5.1.
@@ 3.2 Impact of Input Query Formulation
As illustrated in Section˜2, the input query consists of two components: a text-based graph representation and a question prompt. In Section˜3.2.1 and Section˜3.2.2, we discuss how Transformers perform the substructure extraction task based on these inputs.
@@ 3.2.1 Text-Based Graph Representations
We start with the comparison of different text-based graph representation methods. Specifically, we focus on the two basic methods, which are the neighborhood-based AL and the edge-based EL. To ensure controllable input lengths, we conduct a toy experiment using graphs with 4 to 8 nodes, keeping both representations within a 100-token limit. We then vary the number of Transformer layers to extract substructures (e.g., triangle, square, or pentagon) from the graph representations. Experimental details are provided in the Table˜7 in Section˜E.1 and the results are shown in Figure˜9.
@@
The experimental results indicate that both AL and EL formats allow Transformers to extract substructures from text-based graph representations. As the size of the target substructure increases, both formats require more Transformer layers to achieve more than 80% accuracy. As mentioned in Remark˜3.6, the theoretical results hold for both AL and EL formats. The intuition is that both formats can be transformed into the same binary adjacency matrix $A(G)$, which encodes the structure of graph $G$. This matrix is then vectorized as $\mathsf{vec}(A(G))$ for processing within the model, where we provides the details in Lemma˜D.6 in Section˜D.2.
Although we theoretically show that EL and AL are equivalent in representational power, experimental results reveal that EL requires more Transformer layers to achieve comparable performance. This may be because EL inherently requires more tokens to explicitly represent all edges, whereas AL encodes the same information with fewer tokens and additional padding, benefiting from its more compact structure. In practice, for a fully connected graph, the AL needs $2\times|V|^{2}-|V|$ tokens, while the EL needs $3\times(|V|^{2}-|V|)$ tokens. Therefore, for efficiency, we mainly adopt AL in our discussions. Section˜E.3 provides the details of the training efficiency comparison on AL and EL.